OpenStudy (anonymous):

X-5Y=8 2X-7Y=1 solving systems of equations by eimination

OpenStudy (anonymous):

These problems can be done by two ways : substitution nor elimination.Since you asked to use elimination i suggest to get rid of one variable

OpenStudy (anonymous):

Multiply by -2 to the first equation so you will have -2x+10y=16 then you add to the second equation to get rid of the variable X

OpenStudy (anonymous):

-2x+10y=16 + 2x -7y=1 you will have 3y= 17 then divide both side by 3 you will get y = 17/3 then subtitute back to the equation and solve for x